# MIT-CTP/5802

### v2

@@ Appendix A Network and Training Details
We list the hyperparameters of all studied baselines in Table 7. As for the preprocessing, in the case of GAP and L-GATr we standardize the 4-momentum inputs using a common normalization for each component to preserve Lorentz equivariance. For the rest of the baselines we perform ordinary standarization.
@@
Architecture 128 channels 5 layers 128 channels 4 layers 128 channels 8 heads 8 blocks 96 scalar ch. 96 multivector ch. 8 blocks 72 scalar ch. 8 multivector ch. 4 blocks 32 scalar ch. 32 multivector ch. 8 heads 8 blocks
@@
We provide the L-GATr hyperparamters for top tagging without pre-training in Table 8. All inputs are preprocessed with a 20 GeV scale factor. L-GATr is trained by minimizing a binary cross entropy loss on the class labels.
Pre-training and full training on JetClass is performed by training L-GATr on the full 100M events over $10^{6}$ iterations. The L-GATr architecture and training hyperparameters is the same that we used for the ordinary top tagging. The only differences are that we now work with 10 output channels and train on a cross entropy loss to accommodate multiclass training, and we use a batch size of 512 to maximize dataset exposure. As for the inputs, the 4-momenta are again scaled by the 20 GeV scale, and the kinematic functions are standardized following the prescription presented in Ref. [59].
@@
We summarize the architecture and training hyperparameters of all four generator baselines in Table 9. We split each dataset into 98% for training and 1% each for validation and testing.
In the classifier test, we train an MLP classifier using binary cross-entropy to distinguish generated events from true events. The classifier inputs include the complete events in the $x$ representation defined in Eq. (32), augmented by all pairwise $\Delta R$ features, as well as the $x$ representations of the reconstructed particles $t,\bar{t},W^{+},W^{-}$. The classifier network consists of 3 layers with 256 channels each. Training is conducted over 500 epochs with a batch size of 1024, a dropout rate of 0.1, and the Adam optimizer with default parameters. We start with an initial learning rate of 0.0003, reducing it by a factor of 10 if validation loss shows no improvement for 5 consecutive epochs. The dataset comprises the full truth data and 1M generated events, with an 80%/10%/10% split for training, testing, and validation, respectively.
@@
Architecture 336 6 layers 108 channels 6 layers 8 heads 32 scalar ch. 16 multivector ch. 8 heads 6 blocks

### v3

@@ Appendix A Network and Training Details
We list the hyperparameters of all studied baselines in Table 8. Their values are chosen following a scan with the goal of maximizing the performance of each architecture. As for the preprocessing, in the case of GAP and L-GATr we standardize the 4-momentum inputs using a common normalization for each component to preserve Lorentz equivariance. For the rest of the baselines we perform ordinary standarization.
@@
Architecture
128 channels
5 layers
128 channels
4 layers
128 channels
8 heads
8 blocks
96 scalar ch.
96 multivector ch.
8 blocks
72 scalar ch.
8 multivector ch.
4 blocks
32 scalar ch.
32 multivector ch.
8 heads
8 blocks
@@
We provide the L-GATr hyperparamters for top tagging without pre-training in Table 9. We use the Lion optimizer [82] as an upgrade to the Adam optimizer in our effort to maximize performance as much as possible. All inputs are preprocessed with a 20 GeV scale factor. L-GATr is trained by minimizing a binary cross entropy loss on the class labels.
Pre-training and full training on JetClass is performed by training L-GATr on the full 100M events over $10^{6}$ iterations. The L-GATr architecture and training hyperparameters is the same that we used for the ordinary top tagging. The only differences are that we now work with 10 output channels and train on a cross entropy loss to accommodate multiclass training. We also use a batch size of 512 to maximize dataset exposure. As for the inputs, the 4-momenta are again scaled by the 20 GeV scale, and the kinematic functions are standardized following the prescription presented in Ref. [59].
@@
We summarize the architecture and training hyperparameters of all four generator baselines in Table 10. We split each dataset into 98% for training and 1% each for validation and testing. We use small validation and test datasets in order to speed up the computation of the likelihood performance metrics.
In the classifier test, we train an MLP classifier using binary cross-entropy to distinguish generated events from true events. The classifier inputs include the complete events in the $x$ representation defined in Eq. (32), augmented by all pairwise $\Delta R$ features, as well as the $x$ representations of the reconstructed particles $t,\bar{t},W^{+},W^{-}$. The classifier network consists of 3 layers with 256 channels each. Training is conducted over 500 epochs with a batch size of 1024, a dropout rate of 0.1, and the Adam optimizer with default parameters. We start with an initial learning rate of $3\times 10^{-4}$, reducing it by a factor of 10 if validation loss shows no improvement for 5 consecutive epochs. The dataset comprises the full truth data and 1M generated events, with an 80%/10%/10% split for training, testing, and validation, respectively.
@@
Architecture
336 channels
6 layers
108 channels
6 layers
8 heads
32 scalar ch.
16 multivector ch.
8 heads
6 blocks